# Replacing Softmax Similarity with a Sharpened Angular Similarity: Theory and Practice of Scaling to Billion-Context Attention

### v2

@@ 3.2 The Final Algorithm
At a high level, RACE does not approximate the $N\times N$ score matrix (which would remain quadratic). Instead, it sketches the sufficient statistics needed to compute the attention outputs directly, yielding a linear-time approximation. Fig. 1 illustrates this distinction by focusing on $o_{5}$, the output embedding of token 5 after attention. In Softmax Attention, computing $o_{5}$ requires the entire column of the attention matrix to get the weighted combination of vectors. RACE Attention, in contrast, employs LSH-indexed hash functions to assign the $N$ keys and values (Ks and Vs) into $R$ representative bucket summaries. After this assignment, each query (the $Q$) is hashed into buckets under the same LSH scheme. For example, the output for token 5, $o_{5}$, is obtained by mixing the summaries of tokens assigned to the same bucket as the one mapped by $Q_{5}$ under LSH. The mixing is weighted by soft probability values derived from the well-defined collision probability of the LSH mapping function. Averaging across $L$ independent tables, a standard sketching technique, further reduces variance and stabilizes the approximation. A complete step-by-step expansion is provided in Appendix (fig. 6) and Algorithm 1. Furthermore, we also provide an intuitive visualization of how similarities between tokens look like across buckets in Appendix (fig. 7).
@@
We next formalize the RACE Attention mechanism in Algorithm 1. As described in Section 2.2, one final technical hurdle remains: the RACE algorithm is non-differentiable. We get around the non-differentiability of RACE sketches by replacing discrete bucket assignments with soft probabilities and using standard cross-entropy loss, preserving differentiability for end-to-end training. Algorithm 1 consists of three key stages: (i) Soft bucketization: Each query/key $x\in\mathbb{R}^{d}$ is randomly projected via $W^{(\ell)}$ hyperplanes and softly assigned to $R=2^{P}$ corners with distribution $\phi^{(\ell)}(x)$ (steps 2–4), (ii) Bucket aggregation: For each table $\ell$, we form per-bucket statistics by accumulating key weights and their weighted values, namely the mass vector $A^{(\ell)}\in\mathbb{R}^{R}$ and the value-sum matrix $B^{(\ell)}\in\mathbb{R}^{R\times d}$, so that $A^{(\ell)}[r]$ is the total (soft) mass in bucket $r$ and $B^{(\ell)}[r,:]$ is the corresponding sum of values. (step 5), (iii) Global normalization: The algorithm averages across $L$ tables to form $\texttt{Num}=\tfrac{1}{L}\sum_{\ell}\Phi_{Q}^{(\ell)}B^{(\ell)}$ and $\texttt{Den}=\tfrac{1}{L}\sum_{\ell}\Phi_{Q}^{(\ell)}A^{(\ell)}$, and reconstructs the final outputs as $\widehat{O}=\operatorname{diag}(\texttt{Den})^{-1}\,\texttt{Num}$ (steps 7–8).
@@
Computational Complexity: The per-table runtime of Algorithm 1 can be decomposed according to its main steps: Step 2 (random projections) costs $\mathcal{O}(NdP)$, Step 3 (logits over $R=2^{P}$ corners) costs $\mathcal{O}(NPR)$, and Step 5 (bucket aggregation) costs $\mathcal{O}(NRd)$. The global accumulation in Step 7 adds $\mathcal{O}(NRd)$ per table. Thus, the per-table runtime is $\mathcal{O}(NdP+NPR+NRd)=\mathcal{O}(NRd)$, with memory $\mathcal{O}(NR+Rd)$. Across $L$ tables, this becomes $\mathcal{O}(LNRd)$ time and $\mathcal{O}(L(NR+Rd))$ space. Compared to Softmax Attention’s $\mathcal{O}(N^{2}d)$ time and $\mathcal{O}(Nd)$ space (FlashAttention implementation), RACE is more efficient since $R,L\ll N$ and $R,L\ll d$, even for moderate $N$ and $d$.
@@ 3.3 Theoretical Analysis of Algorithm 1
Algorithm 1 is presented in terms of random projections, soft bucketization, and per-bucket aggregation. For analysis it is easier to take a kernel approximation view of attention. Each hash table $\ell=1,\dots,L$ induces a randomized feature map $\phi^{(\ell)}:\mathbb{R}^{d}\to\mathbb{R}^{R}$, where $R=2^{P}$ is the number of hypercube corners, and defines the approximate kernel $\widehat{S}^{(\ell)}_{ij}=\big(\phi^{(\ell)}(Q_{i})\big)^{\top}\phi^{(\ell)}(K_{j}).$ Then, averaging across $L$ independent tables yields $\widehat{S}=\tfrac{1}{L}\sum_{\ell=1}^{L}\widehat{S}^{(\ell)}$. This view places soft RACE Attention in the language of kernel methods: it replaces the angular kernel ($\gamma=P$) in Eq. (4) with the randomized sketch $\widehat{S}$ based on LSH-style features. Since $\phi^{(\ell)}(x)$ is a softmax distribution, the approximate kernel $\widehat{S}$ inherits concentration properties from the underlying random Gaussian projections. This allows us to analyze its deviation from the target angular kernel using standard tools from randomized numerical linear algebra (RandNLA). Our analysis requires the following two mild assumptions on the target kernel $S$:
@@
Several comments are necessary to better understand the above structural conditions. Condition (A1) rules out degenerate cases where a query has vanishing similarity with all keys, which would make the row-normalization in attention unstable. In practice this assumption is mild: with learned representations, attention rows rarely collapse to near-zero mass, so requiring $s_{\min}\geq cN$ simply rules out degenerate cases where a query is effectively isolated (assigns negligible weight to almost all keys). Condition (A2) is even less restrictive: since $S_{ij}\in[0,1]$, the worst case is attained by the all-ones matrix $J_{N}$, whose spectral norm is exactly $\|J_{N}\|_{2}=N$. Thus bounding $\|S\|_{2}\leq C_{2}N$ merely rules out pathological growth beyond this trivial maximum, and is always satisfied up to a constant factor. We are now ready to state our main quality-of-approximation result:
@@
The bound decomposes into a bias term $\mathcal{O}(P/\beta)$ and a variance term $\mathcal{O}(\sqrt{\log(N/\delta)/L})$. Larger $\beta$ reduces the bias, while increasing $L$ reduces the variance. The dependence on $P$ arises because powering the angular kernel by $P$ makes collisions sharper, but soft bucketization (finite $\beta$) smooths out these decisions and introduces additional bias. To keep this bias small, $\beta$ should be scaled with $P$. In particular, as $\beta,L\to\infty$, the approximation error vanishes. In fact, taking $L=\Theta(\log N)$ prevents the variance from exploding. Together, this kernel reinterpretation provides a precise RandNLA lens for analyzing RACE Attention, with $L$, $P$, and $\beta$ jointly governing its accuracy-efficiency trade-offs. The proof of Theorem 2, together with all intermediate lemmas, is deferred to Appendix 6 due to space constraints.
Remark (Causal masking): Our language modeling experiments in Section 4.1 employ RACE Attention with causal masking, implemented efficiently via OpenMP for CPU workloads. See Algorithm 2 in the Appendix for the causal soft RACE Attention algorithm. Our theoretical analysis above applies only to the non-causal setting. Extending the bias–variance guarantees of Theorem 2 and creating an efficient CUDA kernel for the masked case remains an open problem, as the cumulative-sum constraint interacts non-trivially with the random feature construction. A rigorous analysis of causal RACE Attention is an important direction for future work.

### v3

@@ 3.2 The Final Algorithm
At a high level, RACE avoids explicitly approximating the $N{\times}N$ attention matrix, which would otherwise remain quadratic. Instead, RACE sketches the sufficient statistics needed to compute attention outputs directly in linear time. Fig. 1 illustrates this difference using $o_{5}$, the output embedding of token 5. In Softmax Attention, computing $o_{5}$ requires the full attention column, whereas RACE softly assigns queries and keys (Qs and Ks) into $R$ LSH-indexed bucket summaries under the same LSH scheme. The values are weighted by the key probabilities and then combined with the query probabilities. Averaging over $L$ independent hash tables further reduces variance, following standard sketching practice. Comprehensive visualizations are provided in Appendix (figs. 7, 8).
@@
We next formalize the RACE Attention mechanism in Algorithm 1. As described in Section 2.2, one final technical hurdle remains: the RACE algorithm is non-differentiable. We get around the non-differentiability of RACE sketches by replacing discrete bucket assignments with soft probabilities and using standard cross-entropy loss, preserving differentiability for end-to-end training. Algorithm 1 consists of three key stages: (i) Soft bucketization: Each query/key $x\in\mathbb{R}^{d}$ is randomly projected via $W^{(\ell)}$ hyperplanes and softly assigned to $R=2^{P}$ corners with distribution $\phi^{(\ell)}(x)$ (steps 2–4), (ii) Bucket aggregation: For each table $\ell$, we form per-bucket statistics by accumulating key weights and their weighted values, namely the mass vector $A^{(\ell)}\in\mathbb{R}^{R}$ and the value-sum matrix $B^{(\ell)}\in\mathbb{R}^{R\times d}$, so that $A^{(\ell)}[r]$ is the total (soft) mass in bucket $r$ and $B^{(\ell)}[r,:]$ is the corresponding sum of values. (step 5), (iii) Global normalization: The algorithm averages across $L$ tables to form $\texttt{Num}=\tfrac{1}{L}\sum_{\ell}\Phi_{Q}^{(\ell)}B^{(\ell)}$ and $\texttt{Den}=\tfrac{1}{L}\sum_{\ell}\Phi_{Q}^{(\ell)}A^{(\ell)}$, and reconstructs the final outputs as $\widehat{O}=\operatorname{diag}(\texttt{Den})^{-1}\,\texttt{Num}$ (steps 7–8). A useful way to interpret Algorithm 1 is through the kernel perspective introduced in Section 2.2. In classical RACE [10], $P$ random hyperplanes generate a hash $h(x)=\operatorname{sign}(W^{(\ell)}x)$, and two vectors collide under this hash with probability $\Pr[h(Q_{i})=h(K_{j})]=S_{ij}:=\mathrm{sim}(Q_{i},K_{j}),$ which is exactly the $P$-powered angular kernel in Eq. 4 with $\gamma=P$. In soft RACE, we keep this geometric structure intact, but replace the hard sign map with a smooth approximation: we compute a “soft” sign vector $\tanh(W^{(\ell)}x)$ and evaluate its alignment with each of the $R=2^{P}$ corner sign patterns $v_{r}\in\{\pm 1\}^{P}$, assigning $x$ to buckets with nonzero probability via a softmax over these alignments. This turns the discrete collision event of classical RACE into a differentiable quantity while preserving its underlying angular dependence. In particular, vectors with small angular distance still assign most of their mass to the same buckets, mirroring the behavior of the $P$-powered angular kernel. As a result, the per-table quantity $\phi^{(\ell)}(Q_{i})^{\top}\phi^{(\ell)}(K_{j})$ serves as a smooth approximation to the $P$-powered angular similarity that our attention mechanism seeks to approximate. We formalize these kernel quantities in the Section 3.3, where the RACE-based approximation $\widehat{S}_{ij}$ is introduced.
@@
Computational Complexity: The per-table runtime of Algorithm 1 can be decomposed according to its main steps: Step 2 (random projections) costs $\mathcal{O}(NdP)$, Step 3 (logits over $R=2^{P}$ corners) costs $\mathcal{O}(NPR)$, and Step 5 (bucket aggregation) costs $\mathcal{O}(NRd)$. The global accumulation in Step 7 adds $\mathcal{O}(NRd)$ per table. Thus, the per-table runtime is $\mathcal{O}(NdP+NPR+NRd)=\mathcal{O}(NRd)$, with memory $\mathcal{O}(NR+Rd)$. Across $L$ tables, this becomes $\mathcal{O}(LNRd)$ time and $\mathcal{O}(L(NR+Rd))$ space. Compared to FlashAttention-2/3’s $\mathcal{O}(N^{2}d)$ time and $\mathcal{O}(Nd)$ space, RACE is more efficient since $R,L\ll N$ and $R,L\ll d$, even for moderate $N$ and $d$.
@@ 3.3 Theoretical Analysis of Algorithm 1
Algorithm 1 is presented in terms of random projections, soft bucketization, and per-bucket aggregation. For analysis it is easier to take a kernel approximation view of attention. Each hash table $\ell=1,\dots,L$ induces a randomized feature map $\phi^{(\ell)}:\mathbb{R}^{d}\to\mathbb{R}^{R}$, where $R=2^{P}$ is the number of hypercube corners, and defines the approximate kernel $\widehat{S}^{(\ell)}_{ij}=\big(\phi^{(\ell)}(Q_{i})\big)^{\top}\phi^{(\ell)}(K_{j}).$ Then, averaging across $L$ independent tables yields $\widehat{S}=\tfrac{1}{L}\sum_{\ell=1}^{L}\widehat{S}^{(\ell)}$. This view places soft RACE Attention in the language of kernel methods: it replaces the angular kernel ($\gamma=P$) in Eq. 4 with the randomized sketch $\widehat{S}$ based on LSH-style features. Since $\phi^{(\ell)}(x)$ is a softmax distribution, the approximate kernel $\widehat{S}$ inherits concentration properties from the underlying random Gaussian projections. This allows us to analyze its deviation from the target angular kernel using standard tools from randomized numerical linear algebra (RandNLA) [36]. Our analysis requires the following two mild assumptions on the target kernel $S$:
@@
Several comments are necessary to better understand the above structural conditions. Condition (A1) rules out degenerate cases where a query has vanishing similarity with all keys, which would make the row-normalization in attention unstable. In practice this assumption is mild: with learned representations, attention rows rarely collapse to near-zero mass, so requiring $s_{\min}\geq C_{1}N$ simply rules out degenerate cases where a query is effectively isolated (assigns negligible weight to almost all keys). Condition (A2) is even less restrictive: since $S_{ij}\in[0,1]$, the worst case is attained by the all-ones matrix $J_{N}$, whose spectral norm is exactly $\|J_{N}\|_{2}=N$. Thus bounding $\|S\|_{2}\leq C_{2}N$ merely rules out pathological growth beyond this trivial maximum, and is always satisfied with $C_{2}=1$. We are now ready to state our main quality-of-approximation result:
@@
The bound decomposes into a bias term $\mathcal{O}(P/\beta)$ and a variance term $\mathcal{O}(\sqrt{\log(N/\delta)/L})$. Larger $\beta$ reduces the bias, while increasing $L$ reduces the variance. The dependence on $P$ arises because powering the angular kernel by $P$ makes collisions sharper, but soft bucketization (finite $\beta$) smooths out these decisions and introduces an additional bias. To keep this bias small, $\beta$ should be scaled with $P$. In particular, as $\beta,L\to\infty$, the approximation error vanishes. In fact, taking $L=\Theta(\log N)$ prevents the variance from exploding. Together, this kernel reinterpretation provides a precise RandNLA lens for analyzing RACE Attention, with $L$, $P$, and $\beta$ jointly governing its accuracy-efficiency trade-offs. The proof of Theorem 2, together with all intermediate lemmas, is deferred to Appendix 6 due to space constraints.
Remark (Causal masking): Our language modeling experiments (Tables 4 and 8) use causal RACE Attention, implemented efficiently in OpenMP/CUDA (Algorithm 2). Our theory, however, covers only the non-causal setting. Extending Theorem 2’s bias–variance guarantees to the causal case remains an open problem, since the cumulative-sum constraint interacts non-trivially with the random-feature construction. A rigorous causal analysis is a key direction for future work.
@@ 4 Experiments
Table 1: Long-context classification performance on a 40GB A100 GPU. Train/Test denote per-epoch runtimes in seconds, and Acc. denotes accuracy.
Arxiv Long-Document Classification
16K 32K 64K
Method Train $\downarrow$ Test $\downarrow$ Acc. $\uparrow$ Train $\downarrow$ Test $\downarrow$ Acc. $\uparrow$ Train $\downarrow$ Test $\downarrow$ Acc. $\uparrow$
RACE (P=2,L=2) 80.5s 3.9s 70.3% 282s 15s 89.4% 561s 22s 97.14%
RACE (P=3,L=3) 82.4s 4.0s 71.3% 289s 15.6s 90.6% 584s 22.5s 97.92%
RACE (P=4,L=4) 84.7s 4.1s 70.8% 300s 16s 91.1% 594s 22.9s 97.4%
Linear 83.8s 4.0s 67.9% 286s 15.9s 87.3% 591s 22.8s 96.35%
Linformer-128 86s 3.2s 64.1% 296s 10.7s 87.5% 616s 15.2s 97.4%
Performer-256 128s 5.8s 68.9% 449s 24.6s 86.5% 952s 35s 96.61%
FlashAttention2 95.7s 3.7s 69.8% 471s 20s 89.7% 1645s 47s 97%